Stupa Corporation manufactures two products; data are shown below:

Contribution Margin Ratio Relative Sales Mix
Product D 50% 60%
Product F 60% 40%

If Stupa's monthly fixed costs average $200,000, what is its break-even point expressed in sales dollars? (Round the answer to the nearest dollar.)

Multiple Choice

  • $152,632

  • $250,000

  • $370,370

  • $320,000
Answer #1

Weighted average contribution margin=Respective contribution margin*Respective sales mix

=(0.5*0.6)+(0.6*0.4)=0.54

Breakeven=Fixed expenses/Contribution margin ratio

=(200,000/0.54)

=$370370(Approx).